Number of molecules per unit volume of an ideal gas is ___________ (a) PRT (b) PN/RT (c) PT/RN (d) RN/KT

Grade:12th pass

5 Answers

Shivam Vaish
7 years ago
Hmmm nice question this question requires the knowledge of ideal gas equation which is Pv= nrt. So your answer will be n/v= p/rt hence again check your answer.
nainika
7 years ago
i also applied the formula PV=nRT and got the same answer as you did
but the answer given is tghe second option, i.e PN/RT
Ahivam
7 years ago
Yes , by applying PV=nRt (ideal gas equation) , we are getting" n/v =p/RT" , but the answer is pN/RT , which is correct ! Look don`t confuse between "n and N" both are totally different from each other , here "N" represent pressure num
Amit bhirud
6 years ago
No. Of molecules per unit volume means volume=1 and no. Of molecules =no. Of moles × N(Avogadro`s number) .....So, by applying ideal gas equation PV=nRT ,....v=1 .. Therefore, n/1=P/RT ...For molecules multiplying both sides by N...So we get nN/1=PN/RT.. Option (b)
Pranshu Gautam
5 years ago
PV=nRT
(P/RT)=(n/V)
n= no.of mole of gas 
we know that one mole of gas = Avagadro no.
therfore n mole of gas contains nN molecules 
therefore (nN/V) give us no. of molecules per unit volume
on RHS AND LHS multiply by N
therefore ans = PN/RT